OpenStudy (anonymous):

Hi guys, I need help solving my last 2 homework problems. The first one is: INT (dy)/ Squareroot(16 + y^2)?

OpenStudy (anonymous):

what is really easier is to look on the last page of the text, where all the formulas are. it will have \[\int\frac{dx}{\sqrt{a^2+x^2}}=\sinh(\frac{x}{a})\] i am sure

OpenStudy (anonymous):

well it will really say \[\int\frac{dx}{\sqrt{a^2+x^2}}=\sinh^{-1}(\frac{x}{a})\]

OpenStudy (anonymous):

okay, I'm told to solve this by using trigonometric substitution. will the answer differ?

OpenStudy (anonymous):

ick. it will be lots longer that is for sure

OpenStudy (anonymous):

Yeah, it's taking me like 30 minutes to go through one problem, and there's 30 in my homework..my notebook looks more like a math diary now..lol